OpenStudy (luhvon):

What is 2 3/4% in decimal form?

OpenStudy (mariamikayla):

2 3/4%= 11/4%= 11/4/100= 11/4* 1/100= 11/400= 0.0275

OpenStudy (luhvon):

thanks

OpenStudy (mariamikayla):

you're welcome :D